# A Unified Causal Framework for Nonlinear Electrodynamics Black Hole from Courant-Hilbert Approach: Thermodynamics and Singularity

### v1

@@ Abstract
We develop a unified framework for analyzing black hole thermodynamics and spacetime structure in Einstein gravity coupled to causal nonlinear electrodynamics (NED) in asymptotically anti–de Sitter backgrounds. The electromagnetic sector is governed by a Generalized Nonlinear Electrodynamics (GNED) Lagrangian obtained from a root-$T\bar{T}$ deformation constructed via the Courant–Hilbert approach, ensuring both duality invariance and causal propagation. This theory contains ModMax, Generalized Born–Infeld (GBI), and self-dual logarithmic electrodynamics as continuous limits. Within this framework we obtain exact charged AdS black hole solutions and perform a detailed study of their thermodynamic properties, including mass, temperature, entropy, and free energy. The resulting phase structure exhibits van der Waals-type transitions between small and large black holes and features a characteristic swallowtail in the free energy at the critical point.
We further investigate the internal geometry and show that the nature of the central singularity is determined by the matter fields sourcing the spacetime. Analysis of the Kretschmann scalar demonstrates how gravity and electromagnetism jointly control curvature blow-up in charged and ModMax black holes, with gravity providing the leading divergence. In contrast, (GBI) and logarithmic (NED) models imprint distinct singularity profiles governed by the specific functional form of their Lagrangians. Examining the near-origin behavior of the metric allows us to identify the parameter ranges that support an event horizon and to determine when the solution instead becomes a naked singularity. Overall, the comparison across models reveals that different causal NED theories produce identifiable signatures in both the strength of the singularity and the conditions for horizon formation.
@@ 1 Introduction
In references Kubiznak:2012wp; Flores-Alfonso:2020euz, the authors construct the first self-gravitating solutions within the unique, conformally invariant and duality-symmetric ModMax theory of nonlinear electrodynamics Bandos:2020hgy; Bandos:2020jsw, introducing a novel charged Taub-NUT spacetime and various black holes including Reissner-Nordström-type, accelerated, and AdS configurations. A key finding is that the nonlinear parameter acts as a screening factor, allowing for novel extremal black holes where the mass can be smaller than the charge Barrientos:2022bzm. Furthermore, the research develops a Melvin-Bonnor electromagnetic universe and related black holes embedded within it, revealing a novel Kerr-Schild representation for these spacetimes and expanding the spectrum of exact solutions to include vortex-like backgrounds Barrientos:2024umq. These results provide the first examples of accelerated and electromagnetized black holes in nonlinear electrodynamics, opening new avenues for studying strong-field regimes and raising fresh challenges for holographic interpretation Barrientos:2024umq. In Bokulic:2025usc, the authors demonstrate that for a broad class of theories, regular black holes necessitate a mutual dependence between Komar mass and electric/magnetic charges, and prove that electrically charged black holes respecting the Maxwellian weak-field limit cannot have a bounded Kretschmann scalar. A key focus is the conformally invariant ModMax theory, for which they develop generalized Harrison transformations to generate new solutions, including multi-black hole configurations, black diholes, and their dilaton extensions Bokulic:2025ucy. These solutions often exhibit screening effects that alter the charge-to-mass ratio and can produce regular, extremal black holes without conical singularities. Collectively, this work underscores that while sophisticated solution-generating techniques can yield regular and multi-black hole spacetimes, achieving regularity through nonlinear electrodynamics often requires significant theoretical trade-offs Bokulic:2025brf. See more on ModMax black holes in Bokulic:2021dtz; Pantig:2022gih; Kruglov:2021bhs; Kruglov:2022qag; Guzman-Herrera:2023zsv; Zhang:2021qga; EslamPanah:2024gxx; EslamPanah:2024qlu; Ahmed:2025qza; Heidari:2025llu; Canfora:2025gwm.
@@
In this framework, we focus on two recently proposed classes of nonlinear electrodynamic theories. One is generated in closed form by a $T\bar{T}$-deformation, while the other is a perturbative extension up to order $\lambda^{3}$, which reduces to the ModMax theory in the limit $\lambda\to 0$ Babaei-Aghbolagh:2025cni. The GNED model preserves electric–magnetic duality and causality and provides a unified framework that includes, as special limits, various well-known nonlinear models such as ModMax, (GBI) Bandos:2020hgy; Babaei-Aghbolagh:2022itg, self-dual logarithmic electrodynamics Soleng:1995kn; Babaei-Aghbolagh:2025cni, and $q$-deformed theories Russo:2024ptw. These properties make the GNED model a coherent and physically consistent framework for studying the responses of charged black holes and their holographic and thermodynamic behavior in asymptotically AdS spacetimes. In this study, we couple the general Lagrangian $\mathcal{L}_{\mathrm{}}(F)$ to four-dimensional Einstein gravity with a negative cosmological constant $\Lambda\leq 0$. Our primary objective is to compute the electromagnetic potential and thermodynamic quantities of charged black holes and analyze their behavior under nonlinear electrodynamic effects. We focus on electrostatic configurations with spherical, planar, and hyperbolic horizon geometries, where the absence of magnetic fields simplifies the analysis while preserving the essential nonlinear characteristics of the theory. The corresponding bulk action takes the form
@@ 3.5 Causal q-deform Electrodynamics
Finally, causal $q$-deformed Electrodynamics also belongs to the class of models satisfying condition (16), confirming the consistency of the CH-based formalism and its dependence on the nonlinearity parameter $\gamma$. Within the framework of CH representations, a broad class of duality-invariant and causal $q$-deformed nonlinear electrodynamics theories can be systematically constructed by choosing CH-functions with specific fractional powers Babaei-Aghbolagh:2025uoz. The corresponding CH-function takes the following form Russo:2024ptw:
@@ 4.1 $AdS$ black holes for General-Born-Infeld Theory
In Figure (1), we analyze the horizons for the case $k=0$, where the black hole can have either a single horizon or two horizons. A similar structure occurs for $k=1$; that is, for a certain range of parameters, the solution possesses a single horizon, which then transitions to a two-horizon state. The interesting feature arises in the case k=1k=-1, where the two-horizon state always exists, and the single-horizon state does not occur. This is a crucial point in our discussion, as it highlights a fundamental difference between the geometries. The consequences of this difference are reflected in the resulting thermodynamic parameters.
@@ 4.2 $AdS$ black holes for Logarithmic Theory
In this section, we analyze black hole and thermodynamic solutions in logarithmic nonlinear electrodynamics Soleng:1995kn; Kruglov:2019ybs; Ali:2020jeb; Ali:2023zgm, defined by the CH function (23). This model, a causal and duality-preserving transformation of Maxwell’s theory, has recently been investigated in the causal Courant-Hilbert framework Russo:2024xnh, which clarifies its connection to $T\bar{T}$-like deformations Babaei-Aghbolagh:2025cni. A refined analysis of causal and duality-symmetric (NED) theory was carried out by Russo and Townsend Russo:2024xnh, who analyzed the logarithmic model in the general CH representation, where $\gamma$ denotes the deformation parameter controlling the logarithmic branch of the theory (26). For an electrostatic configuration with charge $Q$ and gauge field $A_{\mu}=(A_{t}(r),\,0,\,0,\,0)$, the magnetic field vanishes. Under this condition, the Lorentz variable $P$ is defined to be zero. This definition implies that the $U$ vanishes and $V=S$. Consequently, the Lagrangian of the logarithmic theory can be written in the following form:
@@ 4.4 $AdS$ black holes for (GNED) Theories
The antisymmetric tensor $G_{\mu\nu}$ is derived from the (GNED) Lagrangian in (81), by taking its derivative with respect to the electromagnetic field strength tensor. This leads to a specific expression for the $G_{01}$ tensor in " (GNED)" theory by $A_{\mu}=(A_{t}(r),0,0,0)$. Finally, to determine the electric potential $A_{t}(r)$, one must solve a differential equation that comes from combining two of the previously established relations. This differential equation is as follows:
@@
$f(r)(-\Lambda-\frac{f^{\prime}(r)}{r})-\frac{{f(r)}^{2}}{r^{2}}-\frac{e^{-\gamma}Q^{2}}{r^{4}}f(r)+\lambda\frac{e^{-2\gamma}m_{1}Q^{4}}{2r^{8}}f(r)$
@@ 4.4.1 Thermodynamic of GNED-AdS black holes
5 Kretschmann tensor and Singularities of black holes
In this section, we propose that the structure of this singularity is not arbitrary but is intrinsically determined by the fundamental properties of the matter that formed the black hole. The Kretschmann tensor is defined as the full contraction of the Riemann tensor:
@@ 5.2 Naked singularity of $AdS$-GBI black hole
Studies in papers Chen:2023trn; Guo:2022ghl investigate black hole solutions and naked Born-Infeld singularities within a framework of gravity coupled to Born-Infeld electrodynamics. These works show that when the singularity exhibits a specific, well-ordered structure, it allows photons entering the photon sphere(s) to pass through it. We now proceed to examine the singularity structure of the (GBI) theory. A distinctive feature of this model, as derived from its Lagrangian, is the emergence of an energy–density singularity that scales as as $1/r^{6}$. For the electrically charged black holes considered, the parameter $\gamma$ only appears in the combination $e^{-\gamma/2}Q$. This implies that the physical behavior in the (GBI) and (GNED) theories is identical to the case with $\gamma=0$, as it is governed entirely by a single effective charge $\tilde{Q}=e^{-\gamma/2}Q$. This behavior is characterized by the parameter $\gamma$, which is intrinsic to the non-linear electrodynamics of the GBI action. To understand the gravitational implications, we analyze the metric function $f(r)$ obtained by solving the coupled Einstein-GBI field equations. Our objective is to determine how this $\mathcal{O}(1/r^{6})$ singularity, governed by $\gamma$, manifests itself in the geometry near the origin $r\to 0$ and influences the global causal structure of the spacetime. Using the definition of the Kretschmann scalar in (91) and the function $f(r)$ in (45), we compute the series expansion of Kretschmann tensor at $r=0$ for AdS-GBI as follows:
$\mathcal{K}=12\biggl(m-\frac{4e^{-3/4\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})\Gamma(\tfrac{5}{4})}{3\pi^{1/2}\lambda^{1/4}}\biggr)^{2}r^{-6}+\mathcal{O}(r^{-5}).$ (93)
A spacetime configuration represents a black hole solution if and only if it possesses an event horizon. The critical boundary that distinguishes a black hole from a naked singularity is defined by the existence of an extremal black hole. To identify this separatrix in parameter space, we analyze the extremal case at $\Lambda=0$, which has a degenerate horizon at radius $r_{e}$ and a corresponding mass $M_{e}=2\,\omega_{e}\,\Omega_{k}$. The defining conditions for this extremal configuration are:
@@
From Condition (94), we can find the extremal radius $r_{e}$ and the parameter $\omega_{e}$ as follows:
$r_{e}=\frac{\sqrt{k}}{2}\sqrt{\frac{4e^{-\gamma}Q^{2}}{k^{2}}-\lambda};\quad\omega_{e}=\frac{1}{3}k^{3/2}\sqrt{\frac{4e^{\gamma}Q^{2}}{k^{2}}-\lambda}\,+\frac{8e^{-\gamma}Q^{2}\,{}_{2}F_{1}\bigl(\tfrac{1}{4},\tfrac{1}{2},\tfrac{5}{4},-\frac{16e^{-\gamma}Q^{2}\lambda}{k^{2}(\lambda-\frac{4e^{-\gamma}Q^{2}}{k^{2}})^{2}}\bigr)}{3\sqrt{k}\sqrt{\frac{4e^{-\gamma}Q^{2}}{k^{2}}-\lambda}}$ (95)
@@
$\lim_{r\to 0}\,r\,f(r)=-m+\frac{4e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})\Gamma(\tfrac{5}{4})}{3\pi^{1/2}\lambda^{1/4}}$ (96)
@@
$m\,<\,\frac{4e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})\Gamma(\tfrac{5}{4})}{3\pi^{1/2}\lambda^{1/4}}.$ (97)

### v2

@@ Abstract
We develop a unified framework for analyzing black hole thermodynamics and spacetime structure in Einstein gravity coupled to causal nonlinear electrodynamics (NED) in asymptotically anti–de Sitter backgrounds. The electromagnetic sector is governed by a Generalized Nonlinear Electrodynamics (GNED) Lagrangian obtained from a root-$T\bar{T}$ deformation constructed via the Courant–Hilbert approach, ensuring both duality invariance and causal propagation. This theory contains ModMax, Generalized Born-Infeld (GBI), and self-dual logarithmic electrodynamics as continuous limits. Within this framework we obtain exact charged AdS black hole solutions and perform a detailed study of their thermodynamic properties, including mass, temperature, entropy, and free energy. The resulting phase structure exhibits van der Waals-type transitions between small and large black holes and features a characteristic swallowtail in the free energy at the critical point.
we further investigate the internal geometry, showing that the nature of the central singularity is determined by the matter fields that source the spacetime. An analysis of the Kretschmann scalar reveals how mass and electric charge jointly govern curvature divergence in ModMax black holes. We derive an explicit charge-to-mass bound within a causal logarithmic electrodynamic theory. This extends the finite self-energy property of the point charge beyond the standard Born-Infeld model. This bound cleanly distinguishes black holes from naked singularities, showing that naked singularities occur precisely when the mass parameter is smaller than the electromagnetic self-energy of the point charge. This provides a clear energetic criterion for horizon formation in the absence of a Cauchy horizon.
@@ 1 Introduction
In references Kubiznak:2012wp; Flores-Alfonso:2020euz, the authors construct the first self-gravitating solutions within the unique, conformally invariant and duality-symmetric ModMax theory of nonlinear electrodynamics Bandos:2020hgy; Bandos:2020jsw, introducing a novel charged Taub-NUT spacetime and various black holes including Reissner-Nordström-type, accelerated, and AdS configurations. A key finding is that the nonlinear parameter acts as a screening factor, allowing for novel extremal black holes where the mass can be smaller than the charge Barrientos:2022bzm. Furthermore, the research develops a Melvin-Bonnor electromagnetic universe and related black holes embedded within it, revealing a novel Kerr-Schild representation for these spacetimes and expanding the spectrum of exact solutions to include vortex-like backgrounds Barrientos:2024umq. These results provide the first examples of accelerated and electromagnetized black holes in nonlinear electrodynamics, opening new avenues for studying strong-field regimes and raising fresh challenges for holographic interpretation Barrientos:2024umq. In Bokulic:2025usc, the authors demonstrate that for a broad class of theories, regular black holes necessitate a mutual dependence between Komar mass and electric/magnetic charges, and prove that electrically charged black holes respecting the Maxwellian weak-field limit cannot have a bounded Kretschmann scalar. A key focus is the conformally invariant ModMax theory, for which they develop generalized Harrison transformations to generate new solutions, including multi-black hole configurations, black diholes, and their dilaton extensions Bokulic:2025ucy. These solutions often exhibit screening effects that alter the charge-to-mass ratio and can produce regular, extremal black holes without conical singularities. Collectively, this work underscores that while sophisticated solution-generating techniques can yield regular and multi-black hole spacetimes, achieving regularity through nonlinear electrodynamics often requires significant theoretical trade-offs Bokulic:2025brf. See more on ModMax black holes in Bokulic:2021dtz; Pantig:2022gih; Kruglov:2021bhs; Kruglov:2022qag; Guzman-Herrera:2023zsv; Zhang:2021qga; EslamPanah:2024gxx; EslamPanah:2024qlu; Ahmed:2025qza; Heidari:2025llu; Canfora:2025gwm; Barrientos:2025rde.
@@
In this framework, we focus on two recently proposed classes of nonlinear electrodynamic theories. One is generated in closed form by a $T\bar{T}$-deformation, while the other is a perturbative extension up to order $\lambda^{3}$, which reduces to the ModMax theory in the limit $\lambda\to 0$ Babaei-Aghbolagh:2025cni. The GNED model preserves electric–magnetic duality and causality and provides a unified framework that includes, as special limits, various well-known nonlinear models such as ModMax, (GBI) Bandos:2020hgy; Babaei-Aghbolagh:2022itg, causal self-dual logarithmic electrodynamics Russo:2024llm; Babaei-Aghbolagh:2025cni, and $q$-deformed theories Russo:2024ptw. These properties make the GNED model a coherent and physically consistent framework for studying the responses of charged black holes and their holographic and thermodynamic behavior in asymptotically AdS spacetimes. In this study, we couple the general Lagrangian $\mathcal{L}_{\mathrm{}}(F)$ to four-dimensional Einstein gravity with a negative cosmological constant $\Lambda\leq 0$. Our primary objective is to compute the electromagnetic potential and thermodynamic quantities of charged black holes and analyze their behavior under nonlinear electrodynamic effects. We focus on electrostatic configurations with spherical, planar, and hyperbolic horizon geometries, where the absence of magnetic fields simplifies the analysis while preserving the essential nonlinear characteristics of the theory. The corresponding bulk action takes the form
@@ 3.5 Causal q-deform Electrodynamics
Finally, causal $q$-deformed Electrodynamics also belongs to the class of models satisfying condition (16), confirming the consistency of the CH-based formalism and its dependence on the nonlinearity parameter $\gamma$. Within the framework of CH representations, a broad class of duality-invariant and causal $q$-deformed nonlinear electrodynamics theories can be systematically constructed by choosing CH-functions with specific fractional powers Babaei-Aghbolagh:2025uoz. The corresponding CH-function takes the following form Russo:2024ptw; Russo:2024xnh:
@@ 4.1 $AdS$ black holes for General-Born-Infeld Theory
In Fig. (1), we analyze the horizon structure for the case $k=0$, where the black hole may exhibit either a single horizon or two distinct horizons, depending on the values of the parameters. A similar qualitative behavior is observed for $k=1$: within a certain parameter range, the solution admits a single-horizon configuration, which subsequently transitions to a two-horizon phase. In the presence of two horizons, the inner horizon corresponds to the Cauchy horizon. A qualitatively different behavior arises for k=1k=-1. In this case, the two-horizon configuration is always present, while a single-horizon solution never occurs. Consequently, a Cauchy horizon is unavoidable for k=1k=-1. This feature is of particular importance, as it highlights a fundamental distinction between the singularity structure and the interior geometry of black holes arising in nonlinear electrodynamics (NED). The interior geometry of the black hole is highly sensitive to variations in the effective electric charge $(\tilde{Q}=e^{-\gamma/2}Q)$ and coupling parameters $(\lambda)$, indicating that the internal behavior is governed by the competition between the gravitational mass and the self-energy.
As shown in Ref. Hale:2025ezt; Hale:2025urg, when the gravitational mass dominates over the self-energy, the interior geometry resembles that of a Schwarzschild black hole, without Cauchy horizon. In contrast, when the self-energy exceeds the mass, the spacetime develops a Reissner-Nordström–like structure, characterized by the presence of a Cauchy horizon and a timelike singularity. A general form of the black factor $f(r)$ for all electrodynamic theories coupled to gravity can be expressed as
$f(r)=k-\frac{m}{r}+\frac{U(r)}{r}\,$, (48)
where $U(r)$ is the self-energy density of the electromagnetic theory, obtained from the stress-energy tensor $U(r)=-\frac{1}{2\pi}\int T^{0}\,_{0}\,dV$. For any theory, the self-energy density can be expanded near $r=0$ in the form
$U(r)=\frac{U^{(-1)}}{r}+U^{(0)}+rU^{(1)}+\mathcal{O}(r^{2})\,.$ (49)
As shown in Hale:2025ezt, for any theory that, like Maxwell’s theory, has a nonzero (and positive) inverse function $U^{(-1)}$, Cauchy horizons will appear. In the (GBI) case, the expansion of the metric function $f(r)$ for large $r=0$ is given by:
$f(r)=k-\frac{m}{r}+\frac{e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{3\pi^{1/2}\lambda^{1/4}}-\frac{2e^{-\tfrac{\gamma}{2}}Q}{\lambda^{1/2}}+\mathcal{O}(r^{2})\,.$ (50)
Since in this theory $U^{(-1)}=0$ and the model does not admit a Cauchy horizon. We know $U^{(0)}$ denote the self-energy of a point charge located at $r=0$. In this model, this quantity satisfies
$U^{(0)}=\frac{e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{3\pi^{1/2}\lambda^{1/4}},$ (51)
which is finite. Consequently, the self-energy of a charged particle in the theory is regular. [^2] We would like to thank Robie A. Hennigar and Jorge G. Russo for discussions on this point. In Sec. (5), we investigate the internal structure of this class of black holes and determine the exact charge-to-mass ratio that separates causal black holes from singular configurations.
@@ 4.2 $AdS$ black holes for Logarithmic Theory
A general discussion of logarithmic black holes can be found in Soleng:1995kn; Kruglov:2019ybs; Ali:2020jeb; Ali:2023zgm, where it is shown that such theories do not satisfy the causality principle in the presence of a magnetic field. In this section, we analyze black hole solutions in logarithmic nonlinear electrodynamics Russo:2024llm; Russo:2024xnh, defined by the CH function (23). This model, a causal and duality-preserving transformation of Maxwell’s theory, has recently been investigated in the causal Courant-Hilbert framework Russo:2024xnh, which clarifies its connection to $T\bar{T}$-like deformations Babaei-Aghbolagh:2025cni. A refined analysis of causal and duality-symmetric (NED) theory was carried out by Russo and Townsend Russo:2024xnh, who analyzed the logarithmic model in the general CH representation, where $\gamma$ denotes the deformation parameter controlling the logarithmic branch of the theory (26). For an electrostatic configuration with charge $Q$ and gauge field $A_{\mu}=(A_{t}(r),\,0,\,0,\,0)$, the magnetic field vanishes. Under this condition, the Lorentz variable $P$ is defined to be zero. This definition implies that the $U$ vanishes and $V=S$. Consequently, the Lagrangian of the logarithmic theory can be written in the following form:
@@
In the Logarithmic nonlinear electrodynamics black hole, the expansion of the metric function $f(r)$ for large $r=0$ in (70) is given by:
$f(r)=k-\frac{m}{r}+\frac{2\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{9\pi^{1/2}\lambda^{1/4}}-\frac{2\sqrt{2}e^{-\tfrac{\gamma}{2}}Q}{\lambda^{1/2}}+\mathcal{O}(r^{2})\,.$ (73)
Considering the general case in equations (48) and  (49), we can compare its expansion with the expansion of the metric function from the logarithmic model in (73). This comparison yields the condition $U^{(-1)}=0$, which shows this model does not have a Cauchy horizon. By comparing the next-order terms, we see that the expansion self-energy density of the logarithmic theory at $r=0$ is
$U^{(0)}=\frac{2\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{9\pi^{1/2}\lambda^{1/4}},$ (74)
which corresponds to a finite self-energy. This represents the first demonstration that a theory beyond the Born-Infeld model possesses a finite self-energy for a point charge.
@@ 4.4 $AdS$ black holes for (GNED) Theories
The antisymmetric tensor $\mathcal{G}_{\mu\nu}$ is derived from the (GNED) Lagrangian in (87), by taking its derivative with respect to the electromagnetic field strength tensor. This leads to a specific expression for the $\mathcal{G}_{01}$ tensor in (GNED) theory by $A_{\mu}=(A_{t}(r),0,0,0)$. Finally, to determine the electric potential $A_{t}(r)$, one must solve a differential equation that comes from combining two of the previously established relations. This differential equation is as follows:
@@
$f(r)(-\Lambda-\frac{f^{\prime}(r)}{r})-\frac{{f(r)}^{2}}{r^{2}}-\frac{e^{-\gamma}Q^{2}}{r^{4}}f(r)+\lambda\frac{e^{-2\gamma}m_{1}Q^{4}}{2r^{8}}f(r)$ (91)
@@ 4.4.1 Thermodynamic of GNED-AdS black holes
5 Kretschmann scalar and Singularities of black holes
In this section, we propose that the structure of this singularity is not arbitrary but is intrinsically determined by the fundamental properties of the matter that formed the black hole. The Kretschmann scalar is defined as the full contraction of the Riemann tensor:
@@ 5.2 Naked singularity of $AdS$-GBI black hole
Studies in Chen:2023trn; Guo:2022ghl investigate black hole solutions and naked Born-Infeld singularities within a framework of gravity coupled to Born-Infeld electrodynamics. These works show that when the singularity exhibits a specific, well-ordered structure, it allows photons entering the photon sphere(s) to pass through it. We now proceed to examine the singularity structure of the (GBI) theory. A distinctive feature of this model, as derived from its Lagrangian, is the emergence of an energy–density singularity that scales as as $1/r^{6}$. For the electrically charged black holes considered, the parameter $\gamma$ only appears in the combination $e^{-\gamma/2}Q$. This implies that the physical behavior in the (GBI) and (GNED) theories is identical to the case with $\gamma=0$, as it is governed entirely by a single effective charge $\tilde{Q}=e^{-\gamma/2}Q$. This behavior is characterized by the parameter $\gamma$, which is intrinsic to the non-linear electrodynamics of the GBI action. To understand the gravitational implications, we analyze the metric function $f(r)$ obtained by solving the coupled Einstein-GBI field equations. Our objective is to determine how this $\mathcal{O}(1/r^{6})$ singularity, governed by $\gamma$, manifests itself in the geometry near the origin $r\to 0$ and influences the global causal structure of the spacetime. Using the definition of the Kretschmann scalar in (97) and the function $f(r)$ in (45), we compute the series expansion of Kretschmann scalar at $r=0$ for AdS-GBI as follows:
$\mathcal{K}=12\biggl(m-\frac{e^{-3/4\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{3\pi^{1/2}\lambda^{1/4}}\biggr)^{2}r^{-6}+\mathcal{O}(r^{-5}).$ (99)
A spacetime configuration represents a black hole solution if and only if it possesses an event horizon. The critical boundary that distinguishes a black hole from a naked singularity is defined by the existence of an extremal black hole. To identify this separatrix in parameter space, we analyze the extremal case at $\Lambda=0$, which has a degenerate horizon at radius $r_{e}$ and a corresponding mass $M_{e}=2\,m_{e}\,\Omega_{k}$. The defining conditions for this extremal configuration are:
@@
From Condition (100), we can find the extremal radius $r_{e}$ and the parameter $m_{e}$ as follows:
$r_{e}=\frac{\sqrt{k}}{2}\sqrt{\frac{4e^{-\gamma}Q^{2}}{k^{2}}-\lambda};\quad m_{e}=\frac{1}{3}k^{3/2}\sqrt{\frac{4e^{\gamma}Q^{2}}{k^{2}}-\lambda}\,+\frac{8e^{-\gamma}Q^{2}\,{}_{2}F_{1}\bigl(\tfrac{1}{4},\tfrac{1}{2},\tfrac{5}{4},-\frac{16e^{-\gamma}Q^{2}\lambda}{k^{2}(\lambda-\frac{4e^{-\gamma}Q^{2}}{k^{2}})^{2}}\bigr)}{3\sqrt{k}\sqrt{\frac{4e^{-\gamma}Q^{2}}{k^{2}}-\lambda}}$ (101)
@@
$\lim_{r\to 0}\,r\,f(r)=-m+\frac{e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{3\pi^{1/2}\lambda^{1/4}}$ (102)
@@
$m\,<\,\frac{e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{3\pi^{1/2}\lambda^{1/4}}.$ (103)
From the self-energy in Eq. (51), we see that the condition $m<U^{(0)}$ leads to a naked singularity in the black hole.